Find the quadratic equation with integral coefficient whose roots are 9 +- (plus or minus) i
Answer by Theo(13342)

the roots are 9 +/- i

start with x = 9 + i and x = 9 - i

to get the factors, set these equations equal to 0.

you do that as follows:

x = 9 + i
subtract 9 and subtract i from both sides of the equation to get:
x - 9 - i = 0

x = 9 - i
subtract 9 and subtract i from both sides of the equation to get:
x - 9 + i = 0

your factors are:

(x - 9 - i) and (x - 9 + i)

now you have to multiply these factors together:

use your distributive law to do this.

look at the picture to see the calculations.

$$$

the next picture shows the factoring of the quadratic equation to confirm the solution is good.

$$$

any questions, send me an email.